solve and check -7x-18= -144
Answer by ankor@dixie-net.com(22740)

solve and check -7x-18= -144
get rid of all those negatives, multiply by -1
7x + 18 = 144
7x = 144 - 18
7x = 126
x = 126/7
x = 18
:
:
Check
-7(18) - 18 = -144